I implemented a slow servo for green laser thermal control on c1scx.mdl. Ch6,7 of ADC and ch6 of DAC are assigned for this servo as below;
Ch6 of ADC: PDH error signal
CH7 of ADC: PZT feedback signal
CH6 of DAC: feedback signal to thermal of green laser
Note that old EPICS themal control cable is not hooked anymore.
I made a simple MEDM screen(...medm/c1scx/master/C1SCX_BCX_SLOW.adl) linked from GREEN medm screen (C1GCV.adl) on sitemap.
During this work, I noticed that some of the epics switch is not recovered by autoburt. What I noticed is filter switch of SUSPOS, SUSPIT, SUSYAW, SDSEN, and all coil output for ETMX.
I had no idea to fix them, probably Joe knows. I guess other suspensitons has the same problems.